# Get a Grip: Reconstructing Hand-Object Stable Grasps in Egocentric Videos

### 25 Dec 2023 v1

@@ 6.5 Ablation Studies
What happens outside the stable grasp boundary? In Table 4 we show results when we remove the knowledge of the grasp start-end time. We run this ablation on a random subset of ARCTIC-Grasps dataset of 100 sequences, for tractable processing. We extend the start-end time of the boundary by 0.5s on both sides.
Clearly, all methods show a drop in performance, as some of these frames might not even have the hand in grasp of the object. Dynamic has the least drop as it does not have a constraint to optimise.
Static Dynamic 1-DoF
ADD SCA-ADD ADD SCA-ADD ADD SCA-ADD
Within 45.0 31.4 55.0 26.7 54.0 33.3
Outside 11.0 6.9 46.0 15.1 27.0 12.9
Table 4: Within/Outside grasps on ARCTIC-Grasps Sequences.
Ablation on Refinement Table 5 presents the impact of performing two refinements on our method. The final refinement stage alone leads to a small improvement; when using both refinement stages, this leads to an average increase of 3.6% in IOU and 7.4% in SCA@0.6.
Init Refine HO Refine IOU SCA@0.8 SCA@0.6
54.1 6.4 18.9
54.7 9.2 23.4
57.8 10.8 25.2
Table 5: Impact of refinement stages on EPIC-Grasps.
Figure 6: Failure Case in EPIC-GRASPS due to FrankMocap.
@@ 7 Conclusion
In this paper we build a challenging in-the-wild egocentric videos dataset for hand-object reconstruction. The dataset contains 2D masks as pseudo ground-truths and manual temporal labels of stable grasps. By focusing on stable grasps, we can jointly optimise multiple frames given the knowledge of how objects move relative to the hand within grasp. We evaluate the method’s ability to achieve a stable grasp without 3D ground truth by approximating the 2D IOU of object masks along with measuring the stable contact area over time.
By investigation stable grasps, we hope to encourage more methods to quantitatively evaluate methods in-the-wild. An obvious limitation of our approach is its reliance on the knowledge of the category’s CAD model. Current CAD-agnostic methods struggle in-the-wild – we showcase failures in appendix [51]. However, removing the limitation of CAD model knowledge is a clear future direction.
@@
This appendix is structured as follows: in Sec. A we presents statistics of EPIC-Grasps annotation. Sec. B showcases failure of current CAD-agnostic approach on EPIC-Grasps dataset. Sec. C provides further implementation details.
Appendix A Statistics of EPIC-Grasps annotation
In Table 6, we present average, standard deviation and maximum clip length for each of our categories as well as total number of video clips per category. Note that grasping some categories (e.g. Pan) can be significantly longer in duration than others (e.g. Mug).
length (sec)
@@ Appendix A Statistics of EPIC-Grasps annotation
Appendix B Evaluating CAD-agnostic method Ihoi
In our method, we assume knowledge of the CAD model. We also explored the latest works that attempt hand grasps without such knowledge. In this section, we showcase these models to be unusable for our in-the-wild scenes through three examplese, qualitatively.
@@ Appendix B Evaluating CAD-agnostic method Ihoi
The recent shape reconstruction method proposed in [22] requires careful scan of the underlying in-hand object hence is not suitable for testing in our case. Therefore, we evaluate the state-of-the-art image-based CAD-agnostic method Ihoi [51] on the EPIC-Grasps dataset, using a model trained on HO3D [17] images that contain the similar categories: mug and bottle – i.e. the method is aware of these shapes during training. Figure 7 shows Ihoi is unable to generate plausible object meshes.
Appendix C Further Implementation Details
Physical Heuristics In Eq (8) of the main paper we note our usage of physical attraction and repulsion losses $E_{ins}$ and $E_{clo}$.
The term $E_{ins}$ ensures all object vertices are located inside the contact surface of the hand (Fig 8). For each $v_{o}\in V_{o}$, we locate the nearest vertex in hand contact regions, and compute distance along the surface normal of this hand vertex. Object vertices that penetrate into the contact surface will have negative values. We maximise those negative values, truncating the positive ones:
$E_{ins}^{n}$ $=\sum_{v_{o}\in V_{o}}-1*\min(d_{v},0)$ (13)
@@ Appendix C Further Implementation Details
In addition to $E_{ins}$, which pushes the object out of the penetrating region against the hand, we use a balancing loss $E_{clo}$ which pulls the object to touch these contact regions. We here focus on the eight contact regions showcased in Fig 8. For each contact region with hand vertices $\{v_{h}\}_{C}$, the region-to-object distance is defined as the minimum distance of all $(v_{h},v_{o})$ pairs. We use 5 finger tip regions and minimise the average of these region-to-object distances.
$E_{clo}^{n}$ $=\frac{1}{5}\sum_{C}d(\{v_{h}\}_{C},V_{o})$ (15)
@@
Initialisation of poses on EPIC-Grasps For EPIC-Grasps dataset, we manually set initial object relative poses to the common poses of each category, results in 80 total initial poses for 9 object categories.

### 07 Apr 2024 v2

@@ 5.5 Ablation Studies
Figure 8: Qualitative Results on EPIC-Grasps (3 examples/ category); projected reconstruction results and reconstruction in rotated views. Bottom: failure cases due to wrong hand pose (left), extreme occlusion (mid) and limited views (right).
In this section, we ablate the hand pose estimator and loss functions under the in-the-wild setting, on full EPIC-Grasps dataset. The 1-DoF assumption of the method has been ablated against other variants in Tab. 2.
Ablation on the hand pose estimation method. In Tab. 4, we compare HaMeR against FrankMocap [35], a commonly used hand predictor in other works [18, 4, 31]. HaMeR consistently outperforms the results with FrankMocap on all metrics, which is in line with the higher capacity of HaMeR over FrankMocap. We include a detailed table in appendix.
Method IOU SCA@0.8 SCA@0.6
1-DoF + FrankMocap [35] 55.1 7.2 20.3
1-DoF + HaMeR [32] 59.9 14.1 32.9
Table 4: Ablation on the hand pose estimator.
Table 5: Ablation on the physical terms. We highlight our choice of $\lambda_{2}$ and $\lambda_{3}$ (blue).
$\lambda_{2}\quad$ $\lambda_{3}\quad$ IOU SCA@0.8 SCA@0.6
0.0 0.0 62.5 15.1 27.5
0.1 0.1 59.9 14.1 32.9
1.0 1.0 53.4 12.2 25.3
$\mathcal{C}_{o}^{n}$ IOU SCA@0.8 SCA@0.6
51.0 3.5 20.4
59.9 14.1 32.9
Table 5: Ablation on the physical terms. We highlight our choice of $\lambda_{2}$ and $\lambda_{3}$ (blue).
Table 6: Ablation on occlusion-aware mask $\mathcal{C}_{o}^{n}$.
Ablation on physical terms $E_{push}$ and $E_{pull}$. In Tab. 6, we ablate the physical terms $E_{push}$ and $E_{pull}$ introduced in the loss function Eq. 5. Note that while removing physical terms ($\lambda_{2}=\lambda_{3}=0$) slightly improves IOU and SCA@0.8, it significantly decreases SCA@0.6 by $5.4\%$.
Ablation on Occlusion Mask. In Tab. 6, we show that using $\mathcal{C}_{o}^{n}$ is critical in the optimisation, whereas removing $\mathcal{C}_{o}^{n}$ drastically reduces the results.
5.6 Limitations
Whilst results in-the-wild are very promising, our pipeline relies on HaMeR as a first stage. Despite the robustness incorporated by the multiple-view joint optimisation, our method fails when the predicted hand poses are wrong (see Fig. 8). Our method also struggle with extreme occlusions and ambiguity from limited views. The 1-DoF assumption also relies on the knowledge of stable grasp duration, which is liable to fail outside grasps.
Another limitation of our approach is its reliance on the knowledge of the category’s CAD model. Current CAD-agnostic methods [47] struggle in-the-wild – we showcase failures in appendix. However, removing the limitation of CAD model knowledge is a clear future direction.
@@ 6 Conclusion
In this paper we propose the task of hand-object stable grasp reconstruction. By focusing on stable grasps, we can jointly optimise multiple frames given the knowledge of how objects move relative to the hand within grasp. We further build a challenging in-the-wild egocentric videos dataset for the task. The dataset contains 2D masks as pseudo ground truths and manual temporal labels of stable grasps. We evaluate the method’s ability without 3D ground truth by measuring the 2D IOU of object masks along with measuring the stable contact area over time. We show that the constrained object motion achieves better hand-object reconstructions within the stable grasp.
By investigation stable grasps, we hope to encourage more works to quantitatively evaluate methods in-the-wild.
@@
This appendix is structured as follows: Appendix 0.A presents statistics of EPIC-Grasps annotation. Appendix 0.B presents statistics of ARCTIC-Grasps annotation. Appendix 0.C provides the stable grasp study on ARCTIC and HOI4D dataset. Appendix 0.D showcases failure of current CAD-agnostic approach on EPIC-Grasps dataset. Appendix 0.E provides ablation of the hand pose estimation method. Appendix 0.F provides further implementation details.
Appendix 0.A Statistics of EPIC-Grasps annotation
In Table 7, we present average, standard deviation and maximum clip duration for each of our categories as well as the total number of video clips per category in EPIC-Grasps. Note that grasping some categories (e.g. pan) can be significantly longer in duration than others (e.g. mug).
Appendix 0.B Statistics of ARCTIC-Grasps automatic annotation
We use Eq (2) in the main paper to automatically extract stable grasp duration from the 3D ARCTIC [13] dataset, where in-contact area can be computed from ground truth meshes. We set $\tau$ to 0.5 empirically: we inspect the resulting grasps to ensure the start-end time match visual expectation. As our pipeline assumes rigid object, we exclude segments where the hand is changing the object’s articulation.
In Table 8, we present average, standard deviation and maximum clip duration for each category as well as total number of video clips per category in ARCTIC-Grasps. As ARCTIC sequences do not include actual activities (e.g. a box is opened only to be closed again rather than to retrieve something from inside), the stable grasps are shorter than those in the unscripted functional recordings of EPIC-Grasps.
Duration (sec)
@@ Appendix 0.B Statistics of ARCTIC-Grasps automatic annotation
Duration (sec)
Category avg std. max #-clips
box 1.52 0.75 4.23 138
capsulemachine 1.51 0.69 3.60 95
espressomachine 1.72 0.95 4.73 101
ketchup 1.51 0.80 4.83 106
laptop 1.39 0.74 4.90 144
microwave 1.51 0.81 5.53 112
mixer 1.44 0.68 3.30 122
notebook 1.15 0.44 2.83 151
phone 1.30 0.91 8.07 146
scissors 1.60 0.77 4.37 57
waffleiron 1.21 0.58 3.90 131
All 1.42 0.76 8.07 1303
Table 8: ARCTIC-Grasps sequences statistics
Appendix 0.C Stable Grasp Study on ARCTIC and HOI4D datasets.
Figure 9: Metrics on ARCTIC within/outside stable grasps: Comparing object in-contact area (left), finger pose (middle) and relative object rotation w.r.t. the hand. Normalising all stable grasp duration for direct comparison.
Figure 10: Metrics on HOI4D within/outside stable grasps: Comparing object in-contact area (left), finger pose (middle) and relative object rotation w.r.t. the hand. Normalising all stable grasp duration for direct comparison.
In this section, we analyse metrics related to stable grasps: stable contact area, finger pose variations and object rotation within/outside stable grasps. Note that hand wrist pose changes drastically w.r.t. the camera, as shown in Fig 3 in the main paper.
ARCTIC Dataset. In Fig. 9 (top), we plot stable contact area, finger pose variations and object rotation errors on 1303 automatically extracted sequences from ARCTIC-Grasps. Among these three metrics, the stable contact area (top-left) shows the sharpest change within/outside stable grasps. We demonstrate that both the finger poses and the object relative orientation do change during the stable grasps, although their changes are bigger outside grasps. Note that the top-left and top-right plots in Fig. 9 are the same as those in Fig 4 of the main paper.
Furthermore, we manually annotate 40 stable grasps in ARCTIC dataset, by observing the videos solely without using the 3D ground truth. We plot the three metrics in Fig. 9 (bottom). As expected, Fig. 9 (bottom-left) shows that the stable grasp duration corresponds to a consistent in-contact area for manual annotated stable grasps. We also verify that the rotation approximation error of 1-DoF are smaller than Static for manually obtained annotations. These figures confirm that the automatically extracted stable gasps are a good approximation of manual annotations of stable grasps.
HOI4D Dataset. To showcase our assumptions on stable grasp generalise to other datasets with 3D annotations, we replicate the analysis above, on the publicly available HOI4D dataset [28]. We use the Eq (2) again to obtain 3568 stable grasps sequences from the HOI4D. Note that the noise of the 3D ground truth pose in HOI4D are much higher than ARCTIC, we thus increase the contact distance threshold to 1cm (from 0cm in ARCTIC) and set $\tau$ to a lower value 0.3. As in ARCTIC, we also manually annotate 60 stable grasps in HOI4D dataset, by observing the videos, and compute the same metrics.
In Fig. 10, we compare the three metrics on the stable grasp sequences on HOI4D, for both automatic annotations (top) and manually annotated samples (bottom). As expected, the stable contact area shows sharp drops at the temporal boundary of stable grasps. Similarly, we observed that the object do rotate relative to the hand, while the rotation error off the 1-DoF axis are much smaller.
These figures showcase that:
  • Our definition of the stable contact temporal duration extends to multiple 3D-annotated datasets.
  • On a variety of datasets, subjects and manipulated objects, the finger poses exhibit considerable changes during the stable grasp.
  • On both datasets and all sequences, objects rotate within the stable grasp – the static assumption produces a large error within the stable grasp ($11^{\circ}$ in ARCTIC compared to $12^{\circ}$ in HOI4D).
  • Optimising for the rotation around a latent axis, the error off the 1-DoF assumption is significantly smaller and better corresponds to the stable grasp temporal extent. This conclusion is verified on both datasets and both manual and automatically annotated sequences.
Appendix 0.D Evaluating CAD-agnostic method IHOI
In our method, we assume knowledge of the CAD model. We also explored the recent works that attempt hand grasps without such knowledge. In this section, we showcase these models to be unusable for in-the-wild stable grasp reconstruction through three examples, qualitatively.
@@ Appendix 0.D Evaluating CAD-agnostic method IHOI
The shape reconstruction method proposed in [21] requires careful scan of the underlying in-hand object hence is not suitable for testing in our case. We also note that HOLDNet [12] and DiffHOI [48] do not have released code to be used. Therefore, we evaluate the image-based CAD-agnostic method IHOI [47] on the EPIC-Grasps dataset, using a model trained on HO3D [17] images that contain the same categories: mug and bottle – i.e. the method is aware of these shapes during training.
Fig. 11 shows IHOI is unable to generate plausible object meshes. As fingers are mostly occluded in egocentric views, this makes the challenge of reconstruction significantly higher than side-view reconstructions where fingers are mostly visible. While CAD-agnostic reconstruction is the ultimate goal, these approaches are currently far from being useful for our scenario.
Appendix 0.E Ablation on hand pose estimation method
While we use the recent transformer-based approach for estimating hand-pose, HaMeR [32], we test our method using the FrankMocap baseline for hand pose etimation [35]. This is an important ablation as our baseline uses FrankMocap.
In Tab. 9, we showcase that our method, using FrankMocap [35] for hand pose etimation outperforms the baseline HOMan [18] which also uses FrankMocap. Using the same hand initialisation, we outperform HOMan [18] by a wide margin - SCA@0.6 improves from 3.3 to 20.3. Using HaMeR, the method further improves consistently over all categories. Average SCA@0.6 improves by an additional 12.6% to reach 32.9.
Fig. 12 shows that FrankMocap struggles with the hand orientation, resulting incorrect reconstruction results.
Appendix 0.F Further Implementation Details
Physical Heuristics In Eq (5) of the main paper we note our usage of physical attraction and repulsion losses $E_{push}$ and $E_{pull}$.
The term $E_{push}$ ensures all object vertices are located inside the contact surface of the hand (Fig 13). For each $v_{o}\in V_{o}$, we locate the nearest vertex in hand contact regions, and compute distance along the surface normal of this hand vertex. Object vertices that penetrate into the contact surface will have negative values. We maximise those negative values, truncating the positive ones:
$E_{push}^{n}$ $=\sum_{v_{o}\in V_{o}}-1*\min(d_{v},0)$ (7)
@@ Appendix 0.F Further Implementation Details
In addition to $E_{push}$, which pushes the object out of the penetrating region against the hand, we use a balancing loss $E_{pull}$ which pulls the object to touch these contact regions. We here focus on the eight contact regions showcased in Fig 13. For each contact region with hand vertices $\{v_{h}\}_{C}$, the region-to-object distance is defined as the minimum distance of all $(v_{h},v_{o})$ pairs. We use 5 finger tip regions and minimise the average of these region-to-object distances.
$E_{pull}^{n}$ $=\frac{1}{5}\sum_{C}d(\{v_{h}\}_{C},V_{o})$ (9)
@@
HOMan [18] 1-DoF + FrankMocap 1-DoF + HaMeR
Category IOU SCA@0.8 SCA@0.6 IOU SCA@0.8 SCA@0.6 IOU SCA@0.8 SCA@0.6
bottle 56.7 3.4 5.8 66.7 11.7 36.8 72.1 26.1 51.3
bowl 54.4 1.1 1.6 53.5 5.5 14.4 56.2 11.1 25.7
can 48.3 3.4 5.1 53.1 11.0 20.9 54.1 16.5 26.6
cup 56.9 5.5 7.8 64.2 11.5 32.5 65.9 18.4 46.1
glass 55.4 3.0 4.2 58.8 10.7 25.7 62.5 15.7 37.8
mug 59.4 3.9 6.6 61.1 7.0 28.3 62.6 10.8 38.6
pan 48.3 0.5 1.9 39.7 2.0 6.9 49.1 7.0 20.4
plate 61.1 1.0 1.6 61.0 9.0 21.5 65.2 16.3 34.6
saucepan 51.1 0.4 2.4 51.4 1.8 16.2 56.8 8.5 34.6
All 54.9 1.9 3.3 55.1 7.2 20.3 59.9 14.1 32.9
Table 9: Ablation on the hand pose estimation method. Green shows the best per metric, Yellow shows the second best.
Figure 12: Qualitative comparison between HaMeR and FrankMocap, when used with our method.
@@
Initialisation of poses on EPIC-Grasps. For EPIC-Grasps dataset, we manually set initial object relative poses to the common poses of each category, results in 80 total initial poses for 9 object categories.
Initialisation of poses on ARCTIC-Grasps. For ARCTIC-Stable dataset, the initial rotations are generated by clustering the ground-truth rotations, where clustering is performed via the axis-angle representation of the rotation matrix. We initialise 50 rotations and 1 global translation for each (object, left/right hand) pair.